OpenStudy (anonymous):

(2^-2 + 2^-3 ) ^-1 ... help please. :))

OpenStudy (solomonzelman):

\[a^{-b}=\frac{1}{a^b}\]So,\[2^{-2}=~?\]\[2^{-3}=?\] After you re-write the problem without any negative exponents, find the common denominator and add. ` Hint 4 × 9 = 36 ` Then after adding what's inside the parenthesis, Use the \(\large\color{black}{ \bf a^{-b}=\frac{1}{a^{b}} }\) rule, again.

OpenStudy (anonymous):

Then, what's the answer? @SolomonZelman

OpenStudy (anonymous):

2.66666666667